Calculate the salt to acid ratio ([Aˉ]/[HA]) of a solution of an amino acid at pH 2 and 

pKa of 8.95


1
Expert's answer
2021-05-16T20:01:02-0400

"pH=pKa+log\frac{[A^-]}{[HA]}"


"2=8.95+log\frac{[A^-}{[HA]}"


"-6.95=log\frac{[A^-]}{[HA]}"


"log\frac{[A_]}{[HA]}=10^{-6.95}"